A solution containing 19.8 grams of barium chloride is mixed with a solution containing 22.4 grams of aluminum sulfate. What mass of barium sulfate will form?Which substance in the excess reagent? 3BaCl2 (aq) + Al2(SO4)3 (aq) → 3BaSO4 (s) + 2AlCl3 (aq *


1
Expert's answer
2021-11-03T08:14:43-0400

Mass of aluminum sulfate = 115 g


Molar mass of aluminum sulfate = 342.15 g/mol


Moles of aluminum sulfate = 115 g / 342.15 g/mol = 0.336 mol


Mass of barium chloride = 225 g


Molar mass of barium chloride = 208.23 g/mol


Moles of barium chloride = 225 g / 208.23 g/mol = 1.08 mol


For complete reaction, the required mol ratio of aluminum sulfate : barium chloride = 1 : 3


Here, the available mol ratio of aluminum sulfate : barium chloride = 0.336 : 1.08


= 1 : 3.2

Thus, the available mol of barium chloride is more than the required mol of barium chloride.

Therefore, barium chloride is the excess reactant in the reaction.

This implies that aluminum sulfate is the limiting reactant in the reaction.

Mass of silver nitrate = 1.699 g

Molar mass of silver nitrate = 169.87 g/mol

Moles of silver nitrate = 1.699 g / 169.87 g/mol = 0.01000 mol

Theoretically, 3 mol of silver nitrate forms 3 mol of silver chloride

Therefore, 0.01000 mol of silver nitrate forms 0.01000 mol of silver chloride

Molar mass of silver chloride = 143.32 g/mol

Hence, the theoretical yield of silver chloride = 0.01000 mol x 143.32 g/mol = 1.433 g